OpenStudy (anonymous):

base is 2 meters and height is 12 meters a=(1)(2)bh

OpenStudy (anonymous):

I have to find the area of the triangle

OpenStudy (anonymous):

a=(1)/(2)bh Thats the correct one!

Parth (parthkohli):

Area of a triangle = 1/2 * Base * Height => 1/2 * 2 * 12 => 1 * 12 => 12 m square

OpenStudy (anonymous):

then it is (base* height)/2

OpenStudy (anonymous):

Im confused

OpenStudy (anonymous):

confused on what?

OpenStudy (anonymous):

how you got 12 m square

OpenStudy (anonymous):

because you said: base is 2 meters and height is "12" meters a=(1)(2)bh :))

OpenStudy (anonymous):

oh nevermind, i got it

OpenStudy (anonymous):

I didnt figure in 1/2 of 24